OpenStudy (cenaida):

What is the y-intercept of the line? y = -3x A. -3 B. 0 C. 1 D. 3

OpenStudy (cenaida):

@MathFan12

OpenStudy (jdoe0001):

what is the y-intercept of say .... y = -9x + 25 ?

OpenStudy (cenaida):

25

OpenStudy (jdoe0001):

well, keep in mind that \(\bf y = -3x \implies y = -3x + 0\)

OpenStudy (cenaida):

so its 3 right?

OpenStudy (jdoe0001):

ahemm.... http://www.algebra-class.com/image-files/slope-formula-1.gif <---

OpenStudy (jdoe0001):

y = -3x + 0 ^ ^ slope y-intercept

OpenStudy (cenaida):

so its 0

OpenStudy (jdoe0001):

yeap

OpenStudy (cenaida):

ok thank you sooo much

OpenStudy (cenaida):

can you help me on this one? Write the equation of the line with the given slope, -3/4 and y-intercept, -2. A. y = -(3/4)x - 2 B. y = -(3/4)x + 2 C. y = -2x - (3/4) D. y = -2x + (3/4)